# Generalization bounds for mixing processes via delayed online-to-PAC conversions

### v1

@@ 6 Conclusion
We have developed a general framework for deriving generalization bounds for non-i.i.d. processes under a general mixing assumption, via an extension of the online-to-PAC-conversion framework of Lugosi and Neu, (2023). Among other results, this approach has allowed us to prove PAC-Bayesian generalization bounds for such data in a clean and transparent way, and even study classes of dynamic hypotheses under a simple bounded-memory condition. These results provide a clean and tight alternative to the results of (Alquier and Wintenberger,, 2012; Eringis et al.,, 2022). The generality of our approach further demonstrates the power of the Online-to-PAC scheme of Lugosi and Neu, (2023), and in particular our results provide further evidence that this framework is particularly promising for developing techniques for generalization in non-i.i.d. settings. We hope that flexibility of our framework will find further uses and enables more rapid progress in the area.
Acknowledgments.
The authors thank Yoav Freund for helpful discussions.
@@ A.1 The proof of Theorem 1
A.2 The proof of Lemma 2
Assume $n=Kd$ for simplicity:
$M_{n}$ $=-\frac{1}{n}\sum_{t=1}^{n}\langle P_{t},c_{t}\rangle$
$=\frac{1}{dK}\sum_{i=1}^{d}\sum_{t=1}^{K}\langle-P_{i+d(t-1)},c_{i+d(t-1)}\rangle$
We denote $X_{t}^{(i)}=\langle-P_{i+d(t-1)},c_{i+d(t-1)}\rangle$ and we want to bound in high-probability the term $\frac{1}{K}\sum_{t=1}^{K}X_{t}^{(i)}.$ Let also denote $\mathcal{F}_{t}^{(i)}=\mathcal{F}_{i+d(t-1)}$. Then for $i\in[\![1,d]\!]$, we can write using Chernoff’s technique that for all $\lambda>0$ it holds:
$\mathbb{P}\left(\frac{1}{K}\sum_{t=1}^{K}X_{t}^{(i)}\geq u\right)$ $\leq\frac{\mathbb{E}\left[e^{\frac{\lambda}{K}\sum_{t=1}^{K}X_{t}^{(i)}}\right]}{e^{\lambda u}}$
$\leq\mathbb{E}\left[e^{\frac{\lambda}{K}\sum_{t=1}^{K-1}X_{t}^{(i)}}\mathbb{E}\left[e^{\frac{\lambda}{K}X_{K}^{(i)}}\Big{|}\mathcal{F}_{K-1}^{(i)}\right]\right]e^{-\lambda u}.$
Now remark that:
$\mathbb{E}\left[e^{\frac{\lambda}{K}X_{K}^{(i)}}\Big{|}\mathcal{F}_{K-1}^{(i)}\right]$ $=\mathbb{E}\left[e^{\frac{\lambda}{K}(X_{K}^{(i)}-\mathbb{E}[X_{K}^{(i)}|F_{K-1}^{(i)}])}\Big{|}F_{K-1}^{(i)}\right]e^{\frac{\lambda}{K}\mathbb{E}[X_{K}^{(i)}|F_{K-1}^{(i)}]}.$
If we denote $Z=X_{K}^{(i)}-\mathbb{E}[X_{K}^{(i)}|F_{K-1}^{(i)}]$ then $|Z|\leq 2$ and $\mathbb{E}[Z|F_{K-1}^{(i)}]=0$ so via Hoeffding’s lemma:
$\mathbb{E}[e^{\frac{\lambda}{K}Z}]\leq e^{\frac{\lambda^{2}}{2K^{2}}}.$
Now by construction of the $P_{t}$ and because of Lemma 1 it follows that for all $i$, $\mathbb{E}[X_{K}^{(i)}|F_{K-1}^{(i)}]\leq\phi_{d}$. Repeating the same reasoning for each term of the sum yields:
$\mathbb{P}\left(\frac{1}{K}\sum_{t=1}^{K}X_{t}^{(i)}\geq u\right)\leq e^{\frac{\lambda^{2}}{2K}}e^{\lambda\phi_{d}}e^{-\lambda u}.$
Optimzing with $\lambda=K(u-\phi_{d})$ and taking $\delta=e^{-\frac{K(u-\phi)^{2}}{2}}$ it finally holds for any $\delta>0,$ with probability $1-\frac{\delta}{d}$:
$\frac{1}{K}\sum_{t=1}^{K}X_{t}^{(i)}\leq\phi_{d}+\sqrt{\frac{2\log\left(\frac{d}{\delta}\right)}{K}}\,.$
Thus applying a union bound we have with probability $1-\delta$:
$M_{n}\leq\phi_{d}+\sqrt{\frac{2\log\left(\frac{d}{\delta}\right)}{K}}\,,$
which concludes the proof.∎
A.3 Proof of Proposition 2
@@
where we used Assumption 3 in the first inequality, Assumption 4 in the second one, and Assumption 3 again in the last step. This proves the statement. ∎
@@ B.1 Regret Bound for EWA
$P_{t+1}=\operatorname*{arg\ min}_{P\in\Delta_{\mathcal{W}}}\left\{\langle P,c_{t}\rangle+\frac{1}{\eta}\mathcal{D}_{\mathrm{KL}}\left(P\middle\|P_{t}\right)\right\},$
@@
$\mathrm{Regret}(P^{*})\leq\frac{\mathcal{D}_{\mathrm{KL}}\left(P^{*}\middle\|P_{1}\right)}{\eta}+\frac{\eta}{2}\sum_{t=1}^{n}\left\|c_{t}\right\|_{\infty}^{2}.$
We refer the reader to Appendix A.1 of Lugosi and Neu, (2023) for a complete proof of the result above.
@@ B.2 Regret Bound for FTRL
$h(\lambda P+(1-\lambda)P^{\prime})\leq\lambda h(P)+(1-\lambda)h(P^{\prime})-\frac{\alpha\lambda(1-\lambda)}{2}\left\|P-P^{\prime}\right\|^{2}.$
@@
$\mathrm{Regret}_{n}(P^{*})\leq\frac{h(P^{*})-h(P_{1})}{\eta}+\frac{\eta}{2\alpha}\sum_{t=1}^{n}\left\|c_{t}\right\|_{*}^{2}.$
We refer the reader to Appendix A.3 of Lugosi and Neu, (2023) for a complete proof of the results above.
B.3 Details about the reduction of Weinberger and Ordentlich, (2002)

### v3

@@ 6 Conclusion
We have developed a general framework for deriving generalization bounds for non-i.i.d. processes under a general mixing assumption, extending the online-to-PAC-conversion framework of Lugosi and Neu (2023). Among other results, this approach has allowed us to prove PAC-Bayesian generalization bounds for such data in a clean and transparent way, and even study classes of dynamic hypotheses under a simple bounded-memory condition. We now conclude by mentioning links with the most closely related previous works in the literature.
The PAC-Bayesian bound in Corollary 4.5 is closely related to Theorem 19 of Ralaivola et al. (2010). Their bound has the advantage to upper-bound a nonlinear proxy of the generalization error and may imply faster rates if the training error is close to zero. Conversely, the proof of this result is a rather cumbersome application of the heavy machinery developed in the rest of their paper. In comparison, our proofs are direct and straightforward, and generalize easily to other divergences than the relative entropy, as discussed in Section 4.4.
Our results concerning dynamic hypotheses provide a clean and tight alternative to the results of (Alquier and Wintenberger, 2012; Eringis et al., 2022). Our Assumptions 3 and 4 can be both shown to hold under their conditions, and the dependence of our bounds on the parameters appearing in these assumptions are stated explicitly. This is a significant advantage over the guarantees of Eringis et al. (2022), which are stated without clearly spelling out the dependence of the problem parameters they consider. Finally, their bounds are stated in PAC-Bayesian terms, which is just one instance of the variety of generalization bounds that our framework can recover. It is worth mentioning that there are other models for studying non-stationarity in statistical learning (León and Perron, 2004; Simchowitz et al., 2018). While our techniques may not be directly applicable to these settings, we believe that a similar online-learning-based perspective might prove helpful for these purposes too.
We also wish to comment on the concurrent work of Chatterjee et al. (2024) who recently proposed a very similar framework for extending the Online-to-PAC technique of Lugosi and Neu (2023) to deal with non-i.i.d. data. Their assumptions and techniques are closely related to ours, except for the major difference that their reduction does not involve delays. Instead, Chatterjee et al. (2024) make strong assumptions about both the loss function and the online learning algorithm: the losses are assumed to be Lipschitz and the online learning method to be “stable” in a rather strong sense. These assumptions are inherited from Agarwal and Duchi (2012) who made the same type of assumptions in the context of proving excess risk bounds for online-to-batch conversions in non-i.i.d. settings. Thanks to the use of delays, our framework does not require any such assumptions, and yields strictly tighter bounds. We note that this observation can be easily used to improve the guarantees of Agarwal and Duchi (2012): by introducing delays into their algorithmic reduction, one can remove the stability conditions required for their results.
Given these positive results, we believe that our results further demonstrate the power of the Online-to-PAC framework of Lugosi and Neu (2023), and clearly show that it is particularly promising for developing techniques for generalization in non-i.i.d. settings. We hope that the flexibility of our framework will find further uses and enable more rapid progress in the area.
[…]
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 950180). The authors would like to thank Yoav Freund for valuable and insightful discussions during the preparation of this work, and the anonymous reviewers for their constructive feedback.
@@ Appendix A Omitted proofs
A.2 Proof of lemma 3.10
Assume $n=Kd$ and denote $M_{n}=\sum_{t=1}^{n}\langle P_{t},c_{t}\rangle$. Then
$M_{n}$ $=\sum_{i=1}^{d}\sum_{t=1}^{K}\langle P_{i+d(t-1)},c_{i+d(t-1)}\rangle\,.$
We denote $X_{t}^{(i)}=\langle P_{i+d(t-1)},c_{i+d(t-1)}\rangle$ and we want to bound in high-probability the term $M_{n}=\sum_{i=1}^{d}M_{i}$, where $M_{i}=\sum_{t=1}^{K}X_{t}^{(i)}.$ We notice that $f\mapsto\log\mathbb{E}\left[e^{\lambda f}\right]$ is convex. Therefore, from Jensen’s inequality, for any positive $p_{1},\dots,p_{d}$ such that $\sum_{i=1}^{M}p_{i}=1$, and for any $\lambda>0$, it holds that
$\log\mathbb{E}\left[e^{\lambda M_{n}}\right]\leq\sum_{i=1}^{d}p_{i}\log\mathbb{E}\left[e^{\lambda\frac{M_{i}}{p_{i}}}\right]\,.$
Let us denote $\mathcal{F}_{t}^{(i)}=\mathcal{F}_{i+d(t-1)}$, we have for all $i\in[d]$
$\mathbb{E}\left[e^{\lambda\frac{M_{i}}{p_{i}}}\right]=\mathbb{E}\left[e^{\frac{\lambda}{p_{i}}\sum_{t=1}^{K}X_{t}^{(i)}}\right]=\mathbb{E}\left[e^{\frac{\lambda}{p_{i}}\sum_{t=1}^{K-1}X_{t}^{(i)}}\mathbb{E}\left[e^{\frac{\lambda}{p_{i}}X_{K}^{(i)}}\Big{|}\mathcal{F}_{K-1}^{(i)}\right]\right]\,.$
Now remark that
$\mathbb{E}\left[e^{\frac{\lambda}{p_{i}}X_{K}^{(i)}}\Big{|}\mathcal{F}_{K-1}^{(i)}\right]=\mathbb{E}\left[e^{\frac{\lambda}{p_{i}}(X_{K}^{(i)}-\mathbb{E}[X_{K}^{(i)}|F_{K-1}^{(i)}])}\Big{|}F_{K-1}^{(i)}\right]e^{\frac{\lambda}{p_{i}}\mathbb{E}[X_{K}^{(i)}|F_{K-1}^{(i)}]}\,.$
Denote $Z=X_{K}^{(i)}-\mathbb{E}[X_{K}^{(i)}|F_{K-1}^{(i)}]$. Note that $|Z|\leq 2$ and $\mathbb{E}[Z|F_{K-1}^{(i)}]=0$. By Hoeffding’s inequality we have $\mathbb{E}\left[e^{\frac{\lambda}{p_{i}}X_{K}^{(i)}}\Big{|}\mathcal{F}_{K-1}^{(i)}\right]\leq e^{\frac{\lambda^{2}}{2p_{i}^{2}}+{\frac{\lambda\phi_{d}}{p_{i}}}}$. Repeating this reasoning $K$ times yields
$\mathbb{E}\left[e^{\lambda\frac{M_{i}}{p_{i}}}\right]\leq e^{\frac{\lambda^{2}K}{2p_{i}^{2}}+{\frac{\lambda K\phi_{d}}{p_{i}}}}\,.$
Finally,
$\log\mathbb{E}\left[e^{\lambda M_{n}}\right]\leq\sum_{i=1}^{d}p_{i}\left(\frac{\lambda^{2}K}{2p_{i}^{2}}+\frac{\lambda K\phi_{d}}{p_{i}}\right)$
now taking $p_{i}=\frac{1}{d}$ essentially gives $\log\mathbb{E}\left[e^{\lambda M_{n}}\right]\leq n(\frac{\lambda^{2}d}{2}+\lambda\phi_{d})$. Now, from Chernoff’s inequality
$\mathbb{P}\left(\frac{M_{n}}{n}\geq t\right)\leq\mathbb{E}\left[e^{\lambda\frac{M_{n}}{n}}\right]e^{-\lambda t}\leq e^{\frac{\lambda^{2}d}{2n}+\lambda\phi_{d}-\lambda t}\leq e^{-\frac{n(t-\phi_{d})^{2}}{2d}}\,.$
Setting $t=\phi_{d}+\sqrt{\frac{2d\log(\frac{1}{\delta})}{n}}$ concludes the proof.
A.3 Proof of Proposition 5.2
@@
where we used Assumption 3 in the first inequality, Assumption 4 in the second one, and Assumption 3 again in the last step. This proves the statement.
@@ B.1 Regret Bound for EWA
$P_{t+1}=\operatorname*{arg\ min}_{P\in\Delta_{\mathcal{W}}}\left\{\langle P,c_{t}\rangle+\frac{1}{\eta}\mathcal{D}_{KL}(P||P_{t})\right\},$
@@
$\mathrm{Regret}(P^{*})\leq\frac{\mathcal{D}_{\text{KL}}(P^{*}||P_{1})}{\eta}+\frac{\eta}{2}\sum_{t=1}^{n}||c_{t}||_{\infty}^{2}.$
We refer the reader to Appendix A.1 of Lugosi and Neu (2023) for a complete proof of the result above.
@@ B.2 Regret Bound for FTRL
$h(\lambda P+(1-\lambda)P^{\prime})\leq\lambda h(P)+(1-\lambda)h(P^{\prime})-\frac{\alpha\lambda(1-\lambda)}{2}||P-P^{\prime}||^{2}.$
@@
$\mathrm{Regret}_{n}(P^{*})\leq\frac{h(P^{*})-h(P_{1})}{\eta}+\frac{\eta}{2\alpha}\sum_{t=1}^{n}||c_{t}||_{*}^{2}.$
We refer the reader to Appendix A.3 of Lugosi and Neu (2023) for a complete proof of the results above.
B.3 Details about the reduction of Weinberger and Ordentlich (2002)